# Complexity of Local Search for Euclidean Clustering Problems

### v1

@@ 4 Reduction to Clustering Problems
Armed with the PLS-completeness of Odd Min/Max Bisection/Flip (see Lemma 3.9), we now proceed to prove hardness of the Euclidean clustering problems of interest.
@@ 4.1 $k$-Means
We construct a PLS-reduction from Odd Min Bisection/Flip to $k$-Means/Hartigan–Wong. This is done in three steps (see Figure 1). First, we show PLS-completeness of Densest Cut/Flip. This reduction is based on a NP-hardness reduction from Max Cut to Densest Cut [MatulaS90]. The construction of the proof of our PLS-completeness of Densest Cut/Flip is rather simple (we only add a large set of isolated edges), but the analysis of the correctness is quite technical. Second, we show PLS-completeness of 2-Means/Hartigan–Wong by slightly adapting an NP-hardness reduction of $2$-Means [ADHP09]. Finally, we extend this result to $k$-Means/Hartigan–Wong.
Now, we show PLS-completeness for Densest Cut/Flip. Here, we impose the additional constraint that there are no isolated vertices in the reduced instance. This is a technical condition which is utilized in LABEL:lemma:densestcut_to_hartigan for the PLS-hardness of $k$-Means/Hartigan–Wong.
@@
Figure 3: Schematic overview of the reduction used in the proof of Lemma 4.1. On the left side we have an instance of Odd Max Bisection/Flip and on the right side we have the corresponding instance of Densest Cut/Flip. The edges inside of $G$ together with their weights are not depicted but are identical in both instances. Let $(A,B)$ be the partition corresponding to some locally optimal solution of the Densest Cut/Flip instance. Then, $A$ contains exactly one endpoint of each of the $n^{4}$ isolated edges and $|\,|A\cap V(G)|-|B\cap V(G)|\,|=1$.
Lemma 4.1.
[…]
Sketch of Proof for Lemma 4.1. The construction of our Densest Cut/Flip instance is based on the simple idea of adding $n^{4}$ isolated edges to $G$, where $(G,w)$ is an instance of Odd Max Bisection/Flip; see Figure 3. The weight $M$ of these newly added edges has to be chosen carefully: On the one hand, $M$ has to be larger than the maximum weight assigned by $w$, but on the other hand $M$ cannot be too large to ensure a specific structure of locally optimal solutions.
To show this statement, we exploit the special structure of the  Odd Max Bisection/Flip-instance. More precisely, if both parts of the cut differ in their size by more than one, then we can flip an arbitrary vertex from the larger part to the smaller part and increase the total weight of the resulting partition. This is true since $G$ is complete. Based on these observations, we are then able to show the second step, that is, that for each locally optimal solution $(A,B)$, the number of vertices of $V(G)$ in $A$ and the number of vertices of $V(G)$ in $B$ differ by exactly one.
Proof 4.2 (Proof of Lemma 4.1).
@@
[…]
Construction of an auxiliary instance. Let $\widehat{G}=(V,\widehat{E})$ be an instance of Odd Max Bisection/Flip with weights $\widehat{w}$ on the edges and let $n\coloneqq|V|$. We now construct an instance $G=(V,E)$ of Odd Max Bisection/Flip with weight vector $w$ such that both instances share the same locally optimal solutions and the minimal weight and the maximal weight assigned by $w$ are close to each other. Note that $G$ and $\widehat{G}$ share the same vertex set $V$. We define $G$ to be the complete graph on the vertex set $V$. Let $\widehat{w}_{\max}\coloneqq\max_{e\in\widehat{E}}w(e)$. We define the weights $w\in\mathds{Z}_{\geq 0}^{E}$ as follows: For each edge $uv\in\widehat{E}$, we set $w(uv)\coloneqq n^{9}\cdot\widehat{w}_{\max}+w(uv)$ and for each $uv\in E\setminus\widehat{E}$, we set $w(uv)\coloneqq n^{9}\cdot\widehat{w}_{\max}$. Now, let $w_{\min}\coloneqq\min_{e\in E}w(e)$ and $w_{\max}\coloneqq\max_{e\in E}w(e)$. Note that based on the defined weights, we obtain the following property:
@@
[…]
Correctness of the auxiliary instance. We now show that both $G$ and $\widehat{G}$ share the same locally optimal solutions.
Claim 3.
@@
This implies that a solution $(A^{\prime},B^{\prime})$ for $G$ or $\widehat{G}$ improves over $(A,B)$ with respect to $w$ if and only if $(A^{\prime},B^{\prime})$ improves over $(A,B)$ with respect to $\widehat{w}$. Consequently, $G$ and $\widehat{G}$ share the same locally optimal solutions and the reduction is tight since the transition graphs of both instances are identical.
[…]
Implications. The bounds between $w_{\min}$ and $w_{\max}$ (see Equation 2) are important to ensure that one can always improve over unbalanced partition by at least $w_{\min}$ by flipping only a single vertex to the other part of the partition if the cardinality constraint is omitted.
Claim 4.
@@
[…]
Construction of the Densest Cut/Flip instance. We now construct the instance $G^{\prime}\coloneqq(V^{\prime},E^{\prime})$ of Densest Cut/Flip as follows; see Figure 3. To obtain $G^{\prime}$, we add a matching of size $n^{4}$ with edge set $E_{M}$ to $G$. Hence, $V^{\prime}\setminus V$ denotes the endpoints of the matching edges of $E_{M}$. Finally, we obtain the weights on the edges of $E^{\prime}$ by extending the weight function $w$: We define the weight of each edge of $E^{\prime}\setminus E=E_{M}$ as $M\coloneqq n\cdot w_{\max}$.
[…]
Correctness. Next, we show that we can map each locally optimal solution of $G^{\prime}$ in polynomial time to a locally optimal solution of $G$ (and thus also for $\widehat{G}$, according to Claim 3). To this end, we first show that each locally optimal solution fulfills two helpful properties. First, we show that in a locally optimal solution all edges of $E_{M}$ are contained in $E(A,B)$.
Claim 5.
@@
Moreover, since $|X|<|Y|-1$, Claim 4 implies that $w(X^{\prime},Y^{\prime})>w(X,Y)+w_{\min}$. This implies that
@@
Hence, flipping an arbitrary vertex of $V\cap Y$ from $B$ to $A$ provides a better solution.
As a consequence, Claims 5 and 6 imply that each locally optimal solution $(A,B)$ for $G^{\prime}$ fulfills $E_{M}\subseteq E(A,B)$ and $|\,|A\cap V|-|B\cap V|\,|=1$. Hence, $(A\cap V,B\cap V)$ is a feasible solution for the instance $G$ of Odd Max Bisection/Flip.
[…]
Mapping of locally optimal solutions. We are now ready to show that for each solution $(A,B)$ for $G^{\prime}$ with $E_{M}\subseteq E(A,B)$ and $|\,|A\cap V|-|B\cap V|\,|=1$, $(X\coloneqq A\cap V,Y\coloneqq B\cap V)$ is a locally optimal solution for $G$ if $(A,B)$ is a locally optimal solution for $G^{\prime}$. We show this statement by contraposition, that is, we show that if $(X,Y)$ is not locally optimal, then $(A,B)$ is not locally optimal. Assume without loss of generality that $|X|=|Y|-1$. Since $(X,Y)$ is not a locally optimal solution for $G$, there exists a vertex $v\in Y$ such that flipping $v$ from $Y$ to $X$ provides a better solution, that is, $w(X\cup\{v\},Y\setminus\{v\})-w(X,Y)>0$. Consider the partition $(A^{\prime},B^{\prime})$ with $A^{\prime}\coloneqq A\cup\{v\}$ and $B^{\prime}\coloneqq B\setminus\{v\}$. Since $|B|=|A|+1$, this implies that $|A|\cdot|B|=|A^{\prime}|\cdot|B^{\prime}|$. Hence, the improvement of $(A^{\prime},B^{\prime})$ over $(A,B)$ is:
@@
[…]
Tightness. We obtain tightness of the reduction by setting $\mathcal{R}$ to be the set of all partitions $(A,B)$ of $G^{\prime}$ that fulfill $E_{M}\subseteq E(A,B)$ and $|\,|A|-|B|\,|=1$. Note that by the above, this implies that $\mathcal{R}$ contains all local optima for $G^{\prime}$. Moreover, for each bisection $(X,Y)$ of $V$, $\mathcal{R}$ contains at least one solution $(A,B)$ with $A\cap V=X$ and $B\cap V=Y$. Hence, the first two properties of a tight PLS-reduction follow.
@@
Due to Claim 6, $(A^{\prime},B^{\prime})$ can be improved by flipping any vertex of $V$ from $B^{\prime}$ to $A^{\prime}$. This includes flipping the vertex $v$ back from $B^{\prime}$ to $A^{\prime}$. Consequently $(A^{\prime},B^{\prime})$ is not improving over $(A,B)$. This implies that for each solution of $\mathcal{R}$ all improving flip-neighbor are in $\mathcal{R}$ as well.
Concluding, to prove the third property of a tight PLS-reduction, it suffices to show that for each arc $(s,s^{\prime})$ in the transition graph for $G^{\prime}$ with $s=(A,B)\in\mathcal{R}$ and $s^{\prime}=(A^{\prime},B^{\prime})\in\mathcal{R}$, there is an arc between $(A\cap V,B\cap V)$ and $(A^{\prime}\cap V,B^{\prime}\cap V)$ in the transition graph for $G$. This follows from the same argumentation as the above proof that each solution $(A,B)$ is locally optimal for $G^{\prime}$ if and only if $(A\cap V,B\cap V)$ is locally optimal for $G$. Consequently, the described reduction is tight.
Next, we show that also the closely related Sparsest Cut is PLS-complete under the Flip neighborhood. Note that Densest Cut and Sparsest Cut are both NP-hard [MatulaS90]. Sparsest Cut is studied intensively in terms of approximation algorithms [AroraRV09] and integrality gaps [KaneM13], and is used to reveal the hierarchical community structure of social networks [MannMO08] and in image segmentation [ShiM00].
Corollary 4.3.
@@
Hence, for any two partitions $(X,Y)$ and $(A,B)$, $\frac{w^{\prime}(X,Y)}{|X|\cdot|Y|}<\frac{w^{\prime}(A,B)}{|A|\cdot|B|}$ if and only if $\frac{w(X,Y)}{|X|\cdot|Y|}>\frac{w(A,B)}{|A|\cdot|B|}$. Thus, both problem instances share the same locally optimal solutions, which proves the correctness of the reduction. Tightness follows by observing that the two instances have the same transition graphs.

### v3

@@ 4 Reduction to Clustering Problems
Armed with the PLS-completeness of Odd Min Bisection/Flip (see Lemma 9), we now proceed to prove hardness of the Euclidean clustering problems of interest.
@@ 4.1 $k$-Means
We provide a tight PLS-reduction from Odd Min Bisection/Flip to $k$-Means/Flip. This is done in three steps (see Figure 1). First, we show PLS-completeness of Densest Cut/Flip. The construction of the proof of our PLS-completeness of Densest Cut/Flip is rather simple (we only add a large set of isolated edges), but the analysis of the correctness is quite technical. Second, we show PLS-completeness of 2-Means/Flip by slightly adapting an NP-hardness reduction of $2$-Means [3]. Finally, we extend this result to $k$-Means/Flip.
Now, we show PLS-completeness of Densest Cut/Flip. We impose the additional constraint that there are no isolated vertices in the reduced instance. This is a technical condition which is utilized in Lemma 12 for the PLS-hardness of $k$-Means/Flip.
Lemma 10.
Sketch of Proof for Lemma 10.
Similar to the obstacle to prove Lemma 6, it is not difficult to show hardness for Densest Cut for swapping two vertices: the main technical hurdle arises when considering Flip, the most fundamental neighborhood. The construction of our Densest Cut/Flip instance is based on the simple idea of adding $n^{4}$ isolated edges to $G$, where $(G,w)$ is an instance of Odd Max Bisection/Flip; see Figure 3. The weight $M$ of these newly added edges has to be chosen carefully: On the one hand, $M$ has to be larger than the maximum weight assigned by $w$, but on the other hand $M$ cannot be too large to ensure a specific structure of locally optimal solutions.
To show this statement, we exploit the special structure of the Odd Max Bisection/Flip-instance. More precisely, if both parts of the cut differ in their size by more than one, then we can flip an arbitrary vertex from the larger part to the smaller part and increase the total weight of the resulting partition. This is true since we will ensure that $G$ is a complete graph where an edge set of larger cardinality is always guaranteed to have a larger total weight. Based on these observations, we then show the second step, that is, that for each locally optimal solution $(A,B)$, the number of vertices of $V(G)$ in $A$ and the number of vertices of $V(G)$ in $B$ differ by exactly one.
@@
Figure 3: Schematic overview of the reduction used in the proof of Lemma 10. On the left side we have an instance of Odd Max Bisection/Flip and on the right side we have the corresponding instance of Densest Cut/Flip. The edges inside of $G$ together with their weights are not depicted but are identical in both instances. Let $(A,B)$ be the partition corresponding to some locally optimal solution of the Densest Cut/Flip instance. Then, $A$ contains exactly one endpoint of each of the $n^{4}$ isolated edges and $|\,|A\cap V(G)|-|B\cap V(G)|\,|=1$.
Proof of Lemma 10.
@@
Construction of an auxiliary instance.
Let $\widehat{G}=(V,\widehat{E})$ be an instance of Odd Max Bisection/Flip with weights $\widehat{w}$ on the edges and let $n\coloneqq|V|$. We now construct an instance $G=(V,E)$ of Odd Max Bisection/Flip with weight vector $w$ such that both instances share the same locally optimal solutions and the minimal weight and the maximal weight assigned by $w$ are close to each other. Note that $G$ and $\widehat{G}$ share the same vertex set $V$. We define $G$ to be the complete graph on the vertex set $V$. Let $\widehat{w}_{\max}\coloneqq\max_{e\in\widehat{E}}w(e)$. We define the weights $w\in\mathds{Z}_{\geq 0}^{E}$ as follows: For each edge $uv\in\widehat{E}$, we set $w(uv)\coloneqq n^{9}\cdot\widehat{w}_{\max}+w(uv)$ and for each $uv\in E\setminus\widehat{E}$, we set $w(uv)\coloneqq n^{9}\cdot\widehat{w}_{\max}$. Now, let $w_{\min}\coloneqq\min_{e\in E}w(e)$ and $w_{\max}\coloneqq\max_{e\in E}w(e)$. Note that based on the defined weights, we obtain the following property:
@@
Correctness of the auxiliary instance.
We now show that both $G$ and $\widehat{G}$ share the same locally optimal solutions.
Claim.
@@
This implies that a solution $(A^{\prime},B^{\prime})$ for $G$ or $\widehat{G}$ improves over $(A,B)$ with respect to $w$ if and only if $(A^{\prime},B^{\prime})$ improves over $(A,B)$ with respect to $\widehat{w}$. Consequently, $G$ and $\widehat{G}$ share the same locally optimal solutions and the reduction is tight since the transition graphs of both instances are identical. ∎
Implications.
The bounds between $w_{\min}$ and $w_{\max}$ (see Equation 2) are important to ensure that one can always improve over unbalanced partition by at least $w_{\min}$ by flipping only a single vertex to the other part of the partition if the cardinality constraint is omitted.
Claim.
@@
Construction of the Densest Cut/Flip instance.
We now construct the instance $G^{\prime}\coloneqq(V^{\prime},E^{\prime})$ of Densest Cut/Flip as follows; see Figure 3. To obtain $G^{\prime}$, we add a matching of size $n^{4}$ with edge set $E_{M}$ to $G$. Hence, $V^{\prime}\setminus V$ denotes the endpoints of the matching edges of $E_{M}$. Finally, we obtain the weights on the edges of $E^{\prime}$ by extending the weight function $w$: We define the weight of each edge of $E^{\prime}\setminus E=E_{M}$ as $M\coloneqq n\cdot w_{\max}$.
Correctness.
Next, we show that we can map each locally optimal solution of $G^{\prime}$ in polynomial time to a locally optimal solution of $G$ (and thus also for $\widehat{G}$, according to Claim Claim). To this end, we first show that each locally optimal solution fulfills two helpful properties. First, we show that in a locally optimal solution all edges of $E_{M}$ are contained in $E(A,B)$.
Claim.
@@
Moreover, since $|X|<|Y|-1$, Claim Claim implies that $w(X^{\prime},Y^{\prime})>w(X,Y)+w_{\min}$. This implies that
@@
Hence, flipping an arbitrary vertex of $V\cap Y$ from $B$ to $A$ provides a better solution. ∎
As a consequence, Claims Claim and Claim imply that each locally optimal solution $(A,B)$ for $G^{\prime}$ fulfills $E_{M}\subseteq E(A,B)$ and $|\,|A\cap V|-|B\cap V|\,|=1$. Hence, $(A\cap V,B\cap V)$ is a feasible solution for the instance $G$ of Odd Max Bisection/Flip.
Mapping of locally optimal solutions.
We are now ready to show that for each solution $(A,B)$ for $G^{\prime}$ with $E_{M}\subseteq E(A,B)$ and $|\,|A\cap V|-|B\cap V|\,|=1$, $(X\coloneqq A\cap V,Y\coloneqq B\cap V)$ is a locally optimal solution for $G$ if $(A,B)$ is a locally optimal solution for $G^{\prime}$. We show this statement by contraposition, that is, we show that if $(X,Y)$ is not locally optimal, then $(A,B)$ is not locally optimal. Assume without loss of generality that $|X|=|Y|-1$. Since $(X,Y)$ is not a locally optimal solution for $G$, there exists a vertex $v\in Y$ such that flipping $v$ from $Y$ to $X$ provides a better solution, that is, $w(X\cup\{v\},Y\setminus\{v\})-w(X,Y)>0$. Consider the partition $(A^{\prime},B^{\prime})$ with $A^{\prime}\coloneqq A\cup\{v\}$ and $B^{\prime}\coloneqq B\setminus\{v\}$. Since $|B|=|A|+1$, this implies that $|A|\cdot|B|=|A^{\prime}|\cdot|B^{\prime}|$. Hence, the improvement of $(A^{\prime},B^{\prime})$ over $(A,B)$ is:
@@
Tightness.
We obtain tightness of the reduction by setting $\mathcal{R}$ to be the set of all partitions $(A,B)$ of $G^{\prime}$ that fulfill $E_{M}\subseteq E(A,B)$ and $|\,|A|-|B|\,|=1$. Note that by the above, this implies that $\mathcal{R}$ contains all local optima for $G^{\prime}$. Moreover, for each bisection $(X,Y)$ of $V$, $\mathcal{R}$ contains at least one solution $(A,B)$ with $A\cap V=X$ and $B\cap V=Y$. Hence, the first two properties of a tight PLS-reduction follow.
@@
Due to Claim Claim, $(A^{\prime},B^{\prime})$ can be improved by flipping any vertex of $V$ from $B^{\prime}$ to $A^{\prime}$. This includes flipping the vertex $v$ back from $B^{\prime}$ to $A^{\prime}$. Consequently $(A^{\prime},B^{\prime})$ is not improving over $(A,B)$. This implies that for each solution of $\mathcal{R}$ all improving flip-neighbor are in $\mathcal{R}$ as well.
Concluding, to prove the third property of a tight PLS-reduction, it suffices to show that for each arc $(s,s^{\prime})$ in the transition graph for $G^{\prime}$ with $s=(A,B)\in\mathcal{R}$ and $s^{\prime}=(A^{\prime},B^{\prime})\in\mathcal{R}$, there is an arc between $(A\cap V,B\cap V)$ and $(A^{\prime}\cap V,B^{\prime}\cap V)$ in the transition graph for $G$. This follows from the same argumentation as the above proof that each solution $(A,B)$ is locally optimal for $G^{\prime}$ if and only if $(A\cap V,B\cap V)$ is locally optimal for $G$. Consequently, the described reduction is tight. ∎
Next, we show that also the closely related Sparsest Cut is PLS-complete under the Flip neighborhood. Note that Densest Cut and Sparsest Cut are both NP-hard [36]. Sparsest Cut is studied intensively in terms of approximation algorithms [5] and integrality gaps [26], and is used to reveal the hierarchical community structure of social networks [33] and in image segmentation [45].
Corollary 11.
@@
Hence, for any two partitions $(X,Y)$ and $(A,B)$, $\frac{w^{\prime}(X,Y)}{|X|\cdot|Y|}<\frac{w^{\prime}(A,B)}{|A|\cdot|B|}$ if and only if $\frac{w(X,Y)}{|X|\cdot|Y|}>\frac{w(A,B)}{|A|\cdot|B|}$. Thus, both problem instances share the same locally optimal solutions, which proves the correctness of the reduction. Tightness follows by observing that the two instances have the same transition graphs. ∎